10mg of potassium was deposited when a current of 0.5A was passed through fused potassium chloride for 1.5hr at 25°C. Calculate the electrochemical equivalent of potassium at this temperature.
1
Expert's answer
2020-10-26T14:53:38-0400

When the electrolysis by one Coulomb charge "Q" is performed, the mass of a substance produced at the electrode "m" equals to the electrochemical equivalent of the substance "Z" . According to this definition, the electrochemical equivalent can be calculated as:

"Z = \frac{m}{Q}" .

The charge passed through the system during the electrolysis equals to the product of the current "I" and time "t" (remember that 1A =1 C/s, 1 hr=60·60 s):

"Q = It = 0.5 \text{ A}\cdot 1.5·60·60\text{ s} = 2700" Coulombs (C).

Therefore, the electrochemical equivalent of potassium is (1mg=1·10-3g):

"Z = \frac{10·10^{-3} \text{ g}}{2700 \text{ C}} = 3.70·10^{-6}" g/C.


Answer: the electrochemical equivalent of potassium at 25°C is 3.70·10-6 g/C.